### v2

# Classport: Designing Runtime Dependency Introspection for Java

Abstract
Runtime introspection of dependencies, i.e., the ability to observe which dependencies are currently used during program execution, is fundamental for Software Supply Chain security. Yet, Java has no support for it. We solve this problem with Classport, a blueprint and system that embeds dependency information into Java class files, enabling the retrieval of dependency information at runtime. We evaluate Classport on six real-world projects, demonstrating the feasibility in identifying dependencies at runtime.
keywords:
Software Supply Chain, Java, Maven
1 Introduction
Software systems keep increasing in size and complexity. To speed up implementation and satisfy the growing demand for features, developers often reuse existing libraries [1]. They leverage third-party open-source code, which is integrated into the main application as a dependency. This collection of source code, dependencies, and build tools used to build a final software product is known as the Software Supply Chain [1].
Using third-party dependencies is essential for productivity and reliability, but it also increases the attack surface [2]. Security vulnerabilities that exist in a dependency are inherited by downstream software products, leading to potential compromises. Notable Software Supply Chain incidents include the Log4Shell vulnerability [3], the Codecov compromise [4], and the xz catastrophe [5]. Assessing and hardening the security of the Software Supply Chain requires increasing its so-called “transparency”, i.e., the visibility on its components at all points in time of the software lifecycle [6].
One of the key tools to support transparency is the production of Software Bill of Materials (SBOM[7]. An SBOM is an exhaustive record of the individual components included in a software system [8]. Maintainers can use this list as input for vulnerability assessment to detect malicious or vulnerable software packages. Most of the approaches to SBOM operate statically, by looking for dependencies in metadata files, such as manifests or lock files, only at build time [9, 10, 11]. This has two fundamental limitations. First, it does not reflect the actual dependency usages at runtime, i.e., whether a dependency is actually loaded or used in production [12]. Even if this does not present a direct risk for the application, it makes maintainers and developers waste a lot of time doing unnecessary work by assessing the security of unreachable dependencies, which slows down the response time to the actual security problems. Second, static analysis of dependencies does not support the maintainers or developers to make runtime decisions based on the actually used dependencies. Basically, the fundamental limitation of SBOMs is that all dependency information is lost at runtime.
Runtime dependency introspection solves these limitations. It is the ability of a piece of software to observe which components are executed during a program’s actual run, including component name and component version. Runtime dependency introspection is the necessary building block for advanced security hardening. First, it is crucial for identifying dependencies that are present in the application but never used at runtime. This simplifies vulnerability management by deprioritizing components that are never used and helps in reducing the attack surface through the removal of unused dependencies. Second, runtime dependency introspection is required to be able to restrict execution privileges at runtime, by mapping the accessed resource to the corresponding dependency [13].
The Java runtime is known for its good support of object-oriented introspection (aka class, field, and method reflection) [14]. Yet, it is completely blind to dependencies: it has no native support for dependency introspection at runtime [15]. In Java, the information that uniquely identifies a dependency is present at build time, but it is not preserved at runtime. Heuristics like guessing the origin of dependencies from package names have been implemented in previous works [13]. However, they suffer from fundamental limitations, such as the requirement for dependencies to be compliant with fragile conventions to be correctly identified. Overall, the problem of runtime introspection of dependencies is an open one. To the best of our knowledge, there are currently no principled architecture for identifying which dependencies are actually executed at runtime in Java.
In this paper, we propose Classport, a novel technique for introspecting dependencies at runtime in Java. In essence, Classport brings dependency information to runtime: supplier, dependency name and version (group id, artifact id and version in Maven lingo). Classport operates in two phases by: 1. embedding dependency information into Java binary artifacts through build-time instrumentation, and 2. retrieving this information at runtime through dynamic instrumentation.
We evaluate Classport on six real-world open-source Java projects. Our experimental protocol measures the ability to embed dependency information into the binary artifacts, and assesses the time and disk space overhead caused by the added information. Then, we test Classport’s ability to correctly identify the set of runtime dependencies using the test suite of the project. We also measure the impact on the running application, incl. the runtime overhead.
Our results demonstrate the applicability of Classport to real-world applications. Classport successfully embeds dependency information, accurately identifies runtime dependencies, preserves the functional behavior of the application, and introduces a low overhead.
In summary, the main contributions of this paper are:
  • A novel approach to inspecting Java dependencies at runtime, based on Java annotations. To the best of our knowledge, it is the first ever solution for runtime dependency introspection in Java.
  • Classport, a publicly available prototype implementing the blueprint: Classport embeds information into binary artifacts and provides a runtime agent to retrieve dependency information during execution. [^1] https://github.com/chains-project/classport
  • An evaluation of our novel technique on a set of six real-world applications, demonstrating the feasibility and applicability of our approach. [^2] https://github.com/chains-project/classport-experiments
2 Background
This section provides background on how dependency management works in the Java ecosystem, with a focus on Maven and its limitations regarding runtime visibility of dependencies.
Dependency management by package managers and build systems varies across ecosystems, as they handle resolution, versioning, and transitive dependencies differently. In the Java ecosystem, Maven is one of the available build systems [16]. Within Maven, developers declare dependencies in a file called pom.xml. Each of these dependencies is downloaded as a binary artifact called a JAR [17], and is uniquely identified by Group ID, Artifact ID, and Version [18], this triple being called a GAV coordinate. The Group ID represents the organization or project namespace, e.g., org.apache.commons. The Artifact ID specifies the module or library, e.g., commons-lang3. The Version distinguishes the same library across different releases, e.g., 3.12.0.
At build time, the dependencies of pom.xml are resolved, in order to list the GAV coordinates of both direct and transitive dependencies. This allows Maven to download the exact version to use. In cases of conflicts between multiple versions of the same transitive dependency, Maven’s dependency resolution mechanism selects the version closest to the root of the dependency tree [17]. The version visible in pom.xml is thus not necessarily the same one that will be used at runtime. In Maven, dependency metadata such as the GAV of dependencies is only available at build time; this information is not preserved at runtime at all [15]. The platform makes it impossible to identify the set of dependencies that are actually used at runtime, or to make any runtime decision depending on the dependencies. In other words, Java supports classes, methods, and fields introspection, but not dependency introspection.
In this paper, we address this problem by proposing a blueprint architecture and a prototype for supporting dependency introspection in Java.
3 Design of Classport
Figure 1: Overview of Classport, a novel system that enables runtime dependency introspection in Java.
We propose Classport, a novel technique for runtime dependency introspection in Java. Recall that in Java, dependency metadata is available at build time, but it is absent during execution.
We define runtime dependencies as dependencies that: 1) are fully specified within the binary to be executed, and 2) are available as variables from within the program.
We also provide an open-source implementation of Classport. The main goal of Classport is to embed Software Supply Chain information, i.e., the dependencies, in the final executable artifact of a Java application, and extract this information at runtime, during application execution. Classport achieves this goal through its two key components: the Embedder and the Introspector. The Embedder stores the information about dependencies in Java artifacts, while the Introspector extracts this embedded information with runtime introspection.
Figure 1 presents an overview of Classport and its integration within the Maven lifecycle. Classport processes a target Java Maven project, which contains the source code and the information about build-time dependencies, which are listed in the pom.xml file. The Embedder performs the embed action (step 2), following the Maven compilation phase (step 1). This produces the Java binary artifacts enriched with the dependency information, which are represented in the output box. After the Maven build phase (step 3), the uber-JAR, i.e., the executable archive comprising the project’s dependencies, is run (step 4). During this phase, the Introspector introspects the execution of the Java program and gives as output the runtime dependencies.
The following two subsections detail the technical aspects of the two components of Classport. Next, we present Classport with an end-to-end example in Section 3.3. Finally, we give details about the implementation of Classport, including tools and libraries used, in Section 3.5.
3.1 Embedder
The Embedder is the first key component of Classport, designed to bridge the gap between build-time and runtime visibility of dependency information. In Java Maven projects, dependency metadata is explicitly declared in the pom.xml and fully resolved during the build phase, but this information is not preserved in the final application artifacts. The Embedder addresses this limitation by capturing build-time dependency metadata and embedding it directly into the compiled artifacts, making it accessible at runtime.
The Embedder takes as input the compiled artifacts of a target Maven project, i.e., the project’s class files and the JARs of the project’s dependencies. The compilation phase, step 1 in Figure 1, ensures that all dependencies are resolved and available locally. For each dependency JAR, the Embedder locates the local version that was downloaded and processes each file within the JAR. Depending on the resource it encounters, this processing involves different actions. For class files within a dependency JAR, i.e., compiled Java source files with .class extension, the Embedder embeds dependency metadata, i.e., its group, artifact, and version, using a bytecode transformation. The same process is applied to the application classes, such as P.class in Figure 1, annotated with the project module information. In Figure 1, the embedded information is represented by the @@ symbol.
When the Embedder encounters manifest files with a dependency JAR, such as MANIFEST.MF files, it removes signature-related entries, e.g., digest attributes, and copies the remaining content. This is necessary because modifying the contents of a signed JAR, such as adding metadata to class files, invalidates its signature. However, Classport is designed to be used by application developers and they can resign the JAR with their keys after the embedding process. For instance, consider a project $P$ authored by an application developer $A_{1}$ that depends on a JAR $D$ signed by a third-party library developer $A_{2}$. During embedding, we strip $A_{2}$’s signature from $D$ so that we can modify the classes, and the embedded dependency is then re-signed by $A_{1}$ once the process completes, as in any vendoring process. The resulting uber-JAR is distributed by $A_{1}$, who is responsible for the packaged software. The Embedder ignores signature files and copies all other files without modification. We refrain from altering non-class files. If these other files are used, they must be referenced by other class files — which are themselves embedded. If they are not used, the Java Virtual Machine will not load them, and hence we don’t need to get provenance later. After processing each dependency JAR, it repackages it into a new JAR. It saves it in a directory, which we refer to as the augmented package repository, as shown in Figure 1.
Finally, as shown in step 3 of Figure 1, the augmented package repository, i.e., the one containing the annotated dependencies, is used to build the final uber-JAR with embedded software supply chain information.
The rest of this subsection describes how the Embedder is implemented.
3.1.1 Embedder as a Maven plugin
We use a Maven plugin to implement the Embedder. [^3] https://maven.apache.org/guides/introduction/introduction-to-plugins.html A Maven plugin is a reusable and highly configurable component that adds specific tasks or goals to Maven’s build process. In particular, it is possible to decide in which build phase to execute it by configuring a Maven plugin goal.
Within Classport, we create the custom embed goal. We configure it to fulfill the following needs of the plugin: 1. to have project’s class files available locally. This is fundamental because they have to be embedded. 2. To have dependencies’ class files. Dependencies must be resolved to also have both direct and transitive ones.
The first point is achieved by compiling the target project before executing the plugin. It guarantees that the class files for the project are available for processing. The second point is fulfilled by ensuring that the resolution of dependencies happens before the embedding process [^4] https://maven.apache.org/plugin-tools/apidocs/org/apache/maven/plugins/annotations/ResolutionScope.html. The embedding and then packaging process happens during a single Maven package phase.
3.1.2 Information embedded as Java Annotations
The primary objective of the tool is to introspect dependency information during execution, so it is crucial to embed data in a manner that allows for retrieval at runtime. We use Java Annotations to add such information into the binary class files. We use bytecode transformation for that task. In each class, we create a custom annotation ClassportInfo, from the annotation type shown in Listing LABEL:lst:annotation.
Listing 1: The Classport Java annotation for storing dependency information within the code. The annotation is available at runtime through introspection.
1@Retention(RetentionPolicy.RUNTIME)
2public @interface ClassportInfo {
3 String group(); // G
4 String artefact(); // A
5 String version(); // V
6}
The annotation contains the GAV coordinate, i.e., group, artifact, and version of the dependency to which the class belongs. The Retention annotation [19], @Retention(RetentionPolicy.RUNTIME), is used to make the ClassportInfo annotation visible at runtime. After the injection of Software Supply Chain information by the Embedder, the target application is packaged as a uber-JAR.
3.2 Introspector
Once dependency metadata has been embedded, it has to be retrieved during execution. The Introspector component of Classport enables this by dynamically extracting the embedded dependency metadata at runtime. As illustrated in phase 4 of Figure 1, the Introspector introspects the running application, reading embedded annotations from running classes to recover dependency metadata. This process allows Classport to determine which dependencies are actually used during execution, providing a precise view of the runtime dependency set.
3.2.1 Introspector as Java agent using the Instrumentation API
The Introspector is a Java agent attached to the JVM. The goal of the Introspector is to add code that extracts the dependency annotations. To achieve this, we instrument the methods of the application under study using code transformation. Each time an instrumented method is executed, the extraction logic is triggered, storing the current dependency GAV in a set. To maintain a low overhead, not every method is instrumented. Specifically, since our objective is to have a list of used dependencies, once a dependency GAV is identified and stored in the set, the subsequent methods belonging to that dependency are no longer instrumented. This allows us to drastically reduce the overhead, without modifying the output of Classport. Finally, after execution of the workload, Introspector writes the resulting set of dependency GAVs into a CSV file.
3.3 End-to-end Example
Listing LABEL:lst:diff shows the changes Classport makes on the original bytecode of the class ExtractText in the PDFBox project. [^5] https://github.com/apache/pdfbox In particular, it shows the resulting bytecode modifications after running the Embedder. Classport adds an attribute in the attribute section (lines 12 to 17), which is the representation of the annotation. The keyword RuntimeVisibleAnnotations (line 12) indicates the retention policy that has been set, and that the annotation is visible at runtime. The remaining part of the section represents the key-value pairs of the annotation, i.e., the Group, Version, and Artifact. Finally, the constant pool section (lines 4-9) is consequently updated with the added values.
Listing LABEL:lst:pdfbox-deps shows the CSV output content of the Introspector. It reports the list of dependencies used during the execution of the application. Specifically, the Group, Artifact, and Version are reported for each dependency.
Listing 2: Example of resulting Java bytecode after embedding process.
1Classfile /.../ExtractText.class
2 ...
3 Constant pool:
4 ...
5 #89 = Utf8 SourceFile
6 #90 = Utf8 ExtractText.java
7+ #91 = Utf8 Lio/github/.../ClassportInfo;
8+ #92 = Utf8 group
9+ #93 = Utf8 org.apache.pdfbox
10 ...
11 Attributes:
12+ RuntimeVisibleAnnotations:
13+ @ClassportInfo(
14+ group = "org.apache.pdfbox",
15+ version = "3.0.4",
16+ artefact = "pdfbox-tools",
17+ )
Listing 3: Result of runtime dependency introspection for PDFBox workload execution. Each row is a triple of Group, Artifact, and Version.
1commons-io,commons-io,2.18.0
2commons-logging,commons-logging,1.3.4
3info.picocli,picocli,4.7.6
4org.apache.pdfbox,fontbox,3.0.4
5org.apache.pdfbox,pdfbox,3.0.4
6org.apache.pdfbox,pdfbox-examples,3.0.4
7org.apache.pdfbox,pdfbox-io,3.0.4
8org.apache.pdfbox,pdfbox-tools,3.0.4
9org.apache.pdfbox,preflight,3.0.4
10org.apache.pdfbox,xmpbox,3.0.4
11org.bouncycastle,bcpkix-jdk18on,1.80
12org.bouncycastle,bcprov-jdk18on,1.80
13org.bouncycastle,bcutil-jdk18on,1.80
3.4 Usage
The application developer is expected to run the packaging phase in Maven with our plugin configured. Our Maven plugin automatically embeds the dependency metadata into the compiled classes, producing an enriched JAR that can be published and distributed to package registries. Users of such JARs get the novel ability to introspect the embedded dependency tree at runtime using the Introspector component. Since Classport is automatically done within the packaging phase, keeping the embedded dependency information up-to-date is fully automated in the release process.
3.5 Implementation
Classport is implemented in Java. It is made of two modules: the Embedder uses the ASM [^6] https://asm.ow2.io/ library to manipulate the bytecode in the class files, while the Introspector is a Java agent that uses the Instrumentation API [20]. The implementation is open-source and available on GitHub. [^7] https://github.com/chains-project/classport
4 Experimental Methodology
This section outlines the methodology for our study evaluating the implementation of Classport. First, it presents the research questions that guide our investigation and information about the dataset. Then, it provides a detailed explanation of how we address these research questions. Lastly, it shows the overall experimental setup.
4.1 Research Questions
We evaluate Classport to demonstrate its effectiveness in introspecting runtime dependencies in real-world applications. First, we assess its capability to embed dependency information into Java binary artifacts. Then, we test its behaviour during runtime execution and its ability to retrieve the actual set of running dependencies. Specifically, we investigate the following research questions:
RQ1: To what extent can Classport effectively embed dependencies into Java binary artifacts?
RQ2: To what extent does Classport support runtime inspection of dependencies?
Table 1: Study subjects considered in our experiments.
Project #Module Version Dependencies
PDFBox 13 3.0.4 19
Certificate Ripper 1 2.4.1 5
mcs 1 0.7.3 5
BioJava 14 7.2.4 50
checkstyle 1 10.23.0 37
GraphHopper 11 11.0 172
4.2 Dataset
We aim at evaluating Classport with a dataset of real-world, maintained, open-source Java projects, which use Maven and span a range of application domains, dependency sizes, and usage scenarios. Our dataset includes six projects, with projects varying in complexity, and includes both simple tools such as mcs and large frameworks such as Checkstyle. The most active repository is Checkstyle, which has around 9K stars on GitHub. Table 1 lists each considered Maven project along with its number of modules, version, and number of dependencies. The modules refer to the Maven modules (also called subprojects) of the project. For example, PDFBox has 13 modules and app and core are two of them. The considerred version is the tagged release of the project. The number of dependencies is calculated directly by Maven. We compile and embed all projects for RQ1 and analyze the resulting artifact. For RQ2, we run the Introspector while executing the tests of the projects in the dataset.
4.3 Methodology of RQ1
The first research question aims to evaluate the Embedder. We assess the dependency completeness, class completeness, and the performance overhead of embedding dependency metadata into Java binary artifacts using Classport. This is done by statically analyzing the JAR file produced by Classport.
Dependency Completeness This property ensures that Classport embeds the uber-JAR with dependency metadata of all dependencies. To assess this, we first run the Embedder on the dataset, generating the annotated uber-JARs. We only consider dependencies that are needed for the application’s execution, excluding test-only dependencies, as they do not end up in the uber-JAR. Then, we analyze the uber-JARs to read the annotations inside the binary artifacts. We need to only consider the uber-JAR because it contains all the dependencies (as opposed to a standard JAR which only contains the classes of the project and not the dependencies). We consider the embedding process complete in terms of dependency metadata if the dependency metadata, i.e., each Group ID, Artifact ID, and Version, matches the ground truth retrieved by running the Maven command mvn dependency:list on the project under test.
Class Completeness We assess completeness by verifying that the annotation is embedded in each class file of every dependency in the project. Although the dependency can have files other than class files, such as build manifests and various configuration files [21], we only consider class files for annotation as they are the only ones that can be executed. To verify completeness, we first run the Embedder on the study subjects. Then, we analyse the uber-JAR, looking inside each class file. If all Java class files are annotated with the Classport annotation, the embedding process is complete in terms of class metadata.
Build Time Overhead Build time overhead is measured by comparing the time taken to build the uber-JAR with and without the Embedder. In detail, we run package phase of the Maven build process with and without the Embedder and measure the time taken for each run using the Unix time command. We do the runs 10 times to accommodate for variance and we report the median of the overhead incurred.
Disk Overhead Disk overhead is the additional space required to store the embedded dependency metadata. To measure the disk overhead, we compare the sizes of the uber-JAR before and after embedding. Specifically, we run the same Maven command as in ‘Build Time Overhead’ to build the uber-JAR with and without the Embedder and measure the disk usage for each run using the Unix stat command. We compute the difference in size of the JAR between the two runs and report the percentage of overhead.
4.4 Methodology of RQ2
The second research question evaluates the behavior of Classport at runtime, in particular, its capability to introspect runtime dependencies. The objective is to show that 1) Classport can identify which dependencies are currently running and 2) the target application’s runtime behavior is not affected by Classport.
To evaluate the Introspector, it is necessary to first run the Embedder, to ensure that the embedding process is completed and also that all the dependency metadata is embedded into the uber-JAR. This means that we assume a correct Embedder, as validated by RQ1, and we now assess the Introspector, which is the component dedicated to the runtime introspection.
Runtime Dependency We report the dependencies detected during workload execution. As workload, we run the test suite of each study subject in Table 1 on the embedded project and record the runtime dependencies it triggers. This produces a list of runtime dependencies observed by Classport. The observed list should contains all embedded dependencies which are exercised at runtime.
Overhead To measure the runtime impact of Classport, we measure the time taken to execute the test suite with and without the Classport agent. We also do this measurement 10 times and report the median of the overhead incurred.
Functional correctness For validating that Classport does not break correctness, we run the test suite of the target application and monitor for runtime errors or test failures caused by the agent.
4.5 General Setup
We run the experiments on a server with an Intel i9-10980XE CPU (36 cores), 125 GB RAM, using Java 17.0.15 and Maven 3.8.7. The dataset and experiment scripts are publicly available on GitHub. [^8] https://github.com/chains-project/classport-experiments
5 Experimental Results
This section presents the results of the experiments as detailed in section 4.
5.1 Results for RQ1
Table 2: RQ1 – Static Completeness and Overhead of Embedding Process
Project Completeness Overhead
Dependencies Classes Time (s) Space (MB)
PDFBox-app (app) 12/12 7,914/7,914 11.52%(2.4) 13.82%(1.9)
Certificate-ripper (itself) 4/5 426/426 6.80%(0.2) 13.17%(0.1)
mcs (itself) 4/4 557/557 7.62%(0.3) 9.88%(0.1)
BioJava (biojava-aa-prop) 31/32 7,629/7,629 7.17%(1.2) 17.19%(2.4)
Checkstyle (itself) 34/37 10,691/10,691 11.38%(2.7) 17.20%(3.4)
GraphHopper (graphhoper-web) 122/166 20,965/20,965 16.80%(6.5) 12.96%(6.1)
The results of RQ1 are summarized in Table 2. It presents the effectiveness of Classport in embedding dependency metadata into Java class files across the open-source projects tested. Each row of the table represents the results of a project on our dataset. The first column lists the projects and their executable uberjar from our dataset described in subsection 4.2. The second column reports the completeness of the embedding process considering two aspects. The first aspect is the number of dependencies embedded compared to the total number of dependencies in the project, and the second one is the number of class files containing the Classport annotation, compared to the total number of class files in the original uber-JAR. Then, the table reports the overhead introduced during the build phase by the Embedder. We report the embedding time overhead, i.e., the time added by the embedding process to the build, and the space overhead, i.e., the size of the uber-JAR after embedding. We also show the absolute increase in paranthesis.
Consider the first row of the table, which corresponds to the PDFBox project. Classport successfully embeds all 12 dependencies of PDFBox-app, which is the total number of dependencies required by the application. The output uber-JAR contains 7,914 class files, and all of them are annotated with the dependency metadata. For this project, the embedding process introduces a time overhead of 11.52% and a space overhead of 13.82% in the final uber-JAR.
There are two projects, PDFBox and mcs, which are completely embedded, while the other projects are partially embedded. There are three reasons why an uber-JAR does not have all the dependencies as declared in the pom.xml file. The first one is the case where a dependency JAR does not contain any class files. It may be empty, which happens for dependency com.google.guava:listenablefuture:9999.0-empty-to-avoid-conflict
-with-guava
, or it may contain files such as JavaScript, Kotlin metadata, or other non-class files. Checkstyle and GraphHopper are examples of projects that have dependencies that do not contain any class files.
The second reason is where a dependency is only necessary for compilation and not for runtime, which is called a provided dependency in Maven. This dependency is not included in the uber-JAR by default. Certificate Ripper and Checkstyle are examples of projects that have dependencies that are only necessary for compilation and not for runtime. Finally, the third reason is where classes of a dependency are shadowed by the other. In this case, BioJava has the dependency javax.xml.bind:jaxb-api which shadows jakarta.xml.bind:jakarta.xml.bind-api. Note that shadowing is not a consequence of the Embedder, but a consequence of how Maven packages the uber-JAR [22]. In all three cases, the classes from those dependencies are not executed. Recall that Classport is designed to detect dependencies that are executed during runtime. Therefore, these dependencies are rightfully not detected by Classport.
The embedding process introduces a moderate performance cost during the build phase. On an absolute scale, the embedding time overhead ranges from 0.2s to 6.5s which we claim is acceptable in a build. The space overhead varies from 0.1MB to 6.1MB which is considered acceptable given the introduction of a novel feature: visibility into runtime dependencies. The space overhead depends upon the number of dependencies and identifier names in the core binary content. Certificate Ripper and mcs have about the same number of classes and the exact same number of dependencies embedded, yet the overhead is larger for Certificate Ripper due to longer identifier names.
Answer to RQ1: Our experiments show that Classport completely embeds dependency metadata into Java class files. Our benchmark of six diverse Maven projects demonstrates that Classport works with real-world and complex software. Build time and space overhead are an acceptable trade-off for getting a unique feature absent from the Java stack: runtime visibility into dependencies.
5.2 Results for RQ2
Table 3: RQ2 – List of Runtime Dependencies and Overhead of Introspector
Project Runtime Dependencies Overhead (s) Functional Correctness
PDFBox 13 4.11%(1.9)
Certificate-ripper 4 2.95%(0.1)
mcs 4 5.96%(0.3)
BioJava 23 8.54%(56.8)
Checkstyle 22 5.28%(8.8)
GraphHopper 104 5.22%(11.0)
We now assess to what extent we are able to obtain the dependency information during execution. Table 3 summarizes the effectiveness of Classport in detecting dependencies at runtime across the evaluated applications. The table is divided into three parts: the first part reports the results regarding the detected runtime dependencies, while the second shows the runtime overhead (as opposed to the build overhead in RQ1), and the third one reports the functional correctness.
Detected Runtime Dependencies This is the number of dependencies detected by Classport during the execution of the workload. For Certificate Ripper and mcs, the workload covers the entire set of embedded dependencies. For the other applications, not all embedded dependencies are detected at runtime. On manual inspection, we found that this is due to two reasons. First, some dependencies are not exercised by the workload or are bloated. Such as the dependency io.dropwizard.logback:logback-throttling-appender:1.5.1 of GraphHopper is not used by the workload. Second, some dependencies have classes that are loaded by the Java Virtual Machine (JVM) but not executed. This happens because JVM loads classes to check for type safety, declared exceptions, and extracting metadata from annotations. For example, the dependency com.google.errorprone:error_prone_annotations:2.36.0 only has annotations. Annotations in Java are not executed as they don’t have any methods. For these two reasons, Classport does not detect these dependencies because it only considers dependencies that are executed during runtime.
Overhead. The overhead during execution of ilow for all the study subjects. It varies from 2.95% to 8.54%. The overhead is mostly due to the fact that Classport needs to instrument methods invocation with an agent.
Functional correctness Classport does not break any functionality as the entire test suite runs with the same result with or without introspection. This means that the bytecode instrumentation done by Classport does not introduce bugs on the target application.
Answer to RQ2: Classport can successfully introspect dependencies at runtime, and allows applications to identify which dependencies are being currently executed. Classport does not affect application behavior. The execution overhead is low for all study subjects. To our knowledge, Classport is the first system in the Java ecosystem to support dependency introspection at runtime.
6 Use Cases of Classport
This section describes the possible benefit of having dependency information available at runtime in Java.
6.1 Runtime Permissions per Dependency
Limiting what resources (e.g., network, filesystem) software components can access at runtime is relevant for security. In the context of Software Supply Chain security, a promising strategy is to assign permissions per dependency, ensuring that each third-party library is granted only the privileges it strictly requires. This permission enforcement per dependency requires precise identification of dependencies at runtime [13, 23].
Classport facilitates this approach in Java by retrieving dependency metadata (Group ID, Artifact ID, and Version) at runtime, enabling permission managers to make security decisions based on the dependency’s inherent characteristics.
6.2 Vulnerability Detection at Runtime
Vulnerability detection tools aim to identify known vulnerabilities in third-party libraries. Identifying dependencies that are actually used at runtime is fundamental to avoid false positives or false negatives in vulnerability assessment [24].
Classport addresses this gap by making it possible to observe exactly which dependencies are executed during a program’s run. By embedding GAV metadata into class files and retrieving it dynamically, Classport enables runtime-aware vulnerability detection tools that focus only on actually used components.
7 Related Work
In this section, we position our contribution with respect to related work on embedding dependency metadata into executables and identifying runtime dependencies in Java.
7.1 Embedding Information into Executables and Binaries
Boucher et al. [25] propose a novel approach to identifying software dependencies, introducing ABOM, Automatic Bill Of Material. The main idea of ABOM is to query an executable for its dependencies to facilitate vulnerability detection. ABOM automatically embeds hashes of dependencies into compiled binaries during the build process. These hashes are stored efficiently in a probabilistic data structure called Compressed Bloom Filters. It allows for rapid querying to detect the presence of a specific dependency and hence to identify the vulnerabilities associated with it. However, since the data structure is probabilistic in nature, it can report the presence of a dependency even if it is not actually present. Our approach cannot generate false positives because it directly relies on Maven and a Maven build success implies that all dependencies for the project are resolved.
Seshadi et al. [26] present OmniBOR, which embeds a content-based identifier called gitoid into executables, including Java class files, to link them to an external Artifact Dependency Graph built during compilation. This approach requires build-time instrumentation and external storage. In contrast, our method embeds Maven GAV metadata directly into class files, allowing immediate runtime access via a Java agent without relying on external resolution. Our technique is dedicated to Java and uniquely supports lightweight, runtime introspection of runtime dependencies.
Quatch et al. [27] propose a solution that combines static and dynamic analysis to identify and remove the unused code during the program loading process. Their framework generates function-level dependency graphs and embeds this information into an optional section of the ELF binary. This embedded dependency information is then used by a piece-wise loader at runtime to identify and remove unused code from the program’s memory. Our approach is more general since it brings dependency information from the build-time context to the runtime one and it is not limited to use cases like debloating.
7.2 Runtime Identification of Dependencies
In this section, we report works related to the runtime identification of dependencies.
Ponta et al. [28] present a novel code-centric and usage-based method for detecting, assessing, and mitigating vulnerabilities in open-source software dependencies. This work builds upon their previous approaches [29, 30], and is implemented in the open-source tool Eclipse Steady. [^9] https://github.com/eclipse-steady/steady They use a dynamic call graph of dependencies to check the reachability of the vulnerable code. They instrument the target Java application and search if, during the execution, it reaches the vulnerable code, i.e., if it uses vulnerable methods. The difference with our approach is that they focus on the vulnerable dependencies and not on an overall identification of the actually executed dependencies.
Soto-Valero et al. [31] propose a tool to detect bloated code in Java and to remove it by transforming the bytecode of the compiled project. They use four code coverage tools along with probes to identify methods that are not executed during runtime. If no methods in any class are executed within a dependency, the dependency is considered bloated. They only work at testing time. However, we embed dependency information directly in the class file in JARs, making this information available at production time as well.
Amusuo et al. [13] implement a tool that is able to create a policy file with a mapping of required permissions for every dependency. The policies are then enforced at runtime. To create a policy file, they infer dependency namespaces by assuming that all classes within a JAR share a common root directory path, which is then mapped to the Maven Group and Artifact ID. In contrast, we directly retrieve dependency information at runtime from metadata embedded in the class files, avoiding fragile assumptions based on package name heuristics.
8 Conclusion
In this paper, we proposed Classport, a novel approach to introspect runtime dependencies in Java. Classport embeds dependency information into Java binary artifacts and exposes this information at runtime, during execution. We demonstrated its effectiveness by augmenting six real-world, open-source Maven Java projects with runtime dependency introspection. For all of them, it is possible to fully embed dependency information and to correctly identify the executed dependencies at runtime, with negligible or low overhead.
As future work, we plan to harden Classport against tampering by malicious Java agents. Since dependency metadata is embedded as annotations, it can be modified at runtime by agents that intercept class loading. To address this, we propose to use cryptographic guarantees by signing the class files, including dependency annotations, at build time. Then, one could verify the signature at load time and run time. This would ensure the integrity of dependency metadata as far as runtime.
Acknowledgements
This work was partially supported by the WASP program funded by Knut and Alice Wallenberg Foundation, and by the Swedish Foundation for Strategic Research (SSF). This work was partially supported by the SERICS project (PE00000014) under the MUR National Recovery and Resilience Plan, funded by the European Union - NextGenerationEU. The authors would like to thank Deepika Tiwari for her contributions to the project.